OpenStudy (anonymous):

Which complex number has an absolute value of 5? A.–3 + 4i B.2 + 3i C.7 – 2i D.9 + 4i

OpenStudy (anonymous):

I just replied to get a notification when someone figures it out, but, idk, I would tink it's D, but I am really not sure.

OpenStudy (anonymous):

absolute value of \(a+bi\) is \(\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

so for example the absolute value of \(-3+4i\) is \[\sqrt{3^2+4^2}=\sqrt{9+16}=\sqrt{25}=5\]

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

yw